A normal body temperature is 37 o

C, Kw = 2.4 x 10-14. Calculate [H3O+

] and [OH-

] for 

the body fluid at this temperature.


1
Expert's answer
2020-09-14T08:20:18-0400

The ion product of water Kw=[H3O+]×[OH-].

[H3O+]=[OH-]=√Kw=√2.4×10-14=1.55×10-7.